# Pentagon relation in quantum cluster scattering diagrams

### v4

@@ 5.2. Rank 2 examples
Let us consider the rank 2 case. As mentioned in Example 5.2, all examples with $\delta_{1}\delta_{2}\leq 4$ in Section 3 are positive. Let us consider examples with $\delta_{1}\delta_{2}>4$, namely, QCSDs of nonaffine infinite type. We use the same conventions and notations in Section 3.
@@
(a). $(\delta_{1},\delta_{2})=(2,3)$. This is the case where the nonpositivity of a quantum greedy element is presented in [LLRZ14], and further examined in view of the QCSD in [CFMM20]. We have
@@
Thus, we confirmed again that it is nonpositive in the sense of Definition 5.1. Moreover, in the classical limit, the negative power in the last expression is cancelled by the positive power, and we obtain $\Psi[\mathbf{e}_{1}]^{2}\Psi[(1,1)]^{6}\Psi[\mathbf{e}_{2}]^{3}$. This clarifies the mechanism how the positivity is restored in the classical limit. Thus, we have an answer to Question 2 in Section 5.1. Moreover, comparing the computation in (5.6) with the much more complicated one in [CFMM20, Appendix B], we observe that the pentagon relation significantly simplifies the analysis (just in two lines in (5.6)).
@@ 5.3. Nonpositivity result
Consider a cluster algebra of rank 2 such that its exchange matrix $B$ has the form in (5.3). In [LLRZ14, Conj. 14], it was conjecture that the strong positivity of the quantum greedy basis holds if $\delta_{1}|\delta_{2}$ or $\delta_{2}|\delta_{1}$. The following result is motivated by this conjecture.
@@
We show that under the condition the nonpositivity already emerges at degree 2 just as above Examples (a) and (d). We continue to assume that $\{e_{2},e_{1}\}=1$. By a similar consideration to the examples in Section 5.2, we have, modulo $G^{>2}$,
@@
Note that this is exactly the situation in Example 5.4.
We have the following corollary of Theorem 5.6, where we convert the statement into a necessary condition for the positivity.
@@
Let $\mathfrak{D}^{q}_{\mathfrak{s}}$ be a QCSD of any rank. Then, if $\mathfrak{D}^{q}_{\mathfrak{s}}$ is positive, $\delta_{i}\mid\delta_{j}$ or $\delta_{j}\mid\delta_{i}$ holds for any pair $i\neq j$ such that $b_{ij}\neq 0$ for $B$ in (2.1).
@@
Suppose that $\delta_{i}\nmid\delta_{j}$ holds for some $i$ and $j$ such that $b_{ij}\neq 0$. We consider a QCSD whose wall elements are in the form (4.29). Then, the proof of Theorem 5.6 tells that there is some wall whose wall element is $\Psi_{1/\delta(e_{i}+e_{j}),b}[e_{i}+e_{j}]^{-1}$ for some $b$. Moreover, the wall element $\Psi_{1/\delta(e_{i}+e_{j}),b}[e_{i}+e_{j}]$ arises only from the situation in the proof of Theorem 5.6 due to the construction of a CSD in [GHKK18, Appendix C.1]. Therefore, the wall element $\Psi_{1/\delta(e_{i}+e_{j}),b}[e_{i}+e_{j}]^{-1}$ is never cancelled by other wall elements. ∎
@@ 5.4. Positivity up to degree 4
We are left with the problem to determine whether the converses of Theorem 5.6 and Corollary 5.7 hold or not. Here, we concentrate on the former one and examine the positivity up to some degrees.
@@
(5.13) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\\ &\qquad\times\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},-\frac{6}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},-\frac{4}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},\frac{4}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},\frac{6}{5}}\\ &\qquad\times\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\begin{bmatrix}1\\ 4\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}1\\ 5\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{5},0}.\end{split}$
@@
(5.14) $\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}2\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{3}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-1}\\ &\qquad\times\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},1}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},\frac{3}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-1}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\\ &\qquad\times\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},1}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},-\frac{1}{2}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
@@
(c). Degree 4. The results are similar to the degree 3 case. However, the formulas become lengthy, so that we put them in Appendix B. They also imply the difficulty to prove the positivity in higher degrees by direct computation.
@@ Appendix B Positivity at degree 4
This is a continuation of Example 5.9 for degree 4. The purpose of writing down the explicit expressions is just to demonstrate how the problem becomes complicated even for such a small degree. We have

### v7

@@ 5.2. Rank 2 examples
Let us focus on the rank 2 case. As mentioned in Example 5.5, all examples with $\delta_{1}\delta_{2}\leq 4$ in Section 3 are positive. Let us consider examples with $\delta_{1}\delta_{2}\geq 5$, namely, QCSDs of nonaffine infinite type. We use the same conventions and notations in Section 3.
@@
(a). $(\delta_{1},\delta_{2})=(2,3)$. This is the case where the nonpositivity of a quantum greedy element [LLRZ14] and (quantum) theta function [CFMM24] was shown. We have
@@
Thus, we confirmed again that it is nonpositive in the sense of Definition 5.4. Moreover, in the classical limit, the negative power in the last expression is cancelled by the positive power, and we obtain $\Psi[\mathbf{e}_{1}]^{2}\Psi[(1,1)]^{6}\Psi[\mathbf{e}_{2}]^{3}$. This clarifies the mechanism how the positivity is restored in the classical limit. Thus, we have a clear answer to Question 2 in Section 5.1. Moreover, comparing the computation in (5.5) with the much more complicated one in [CFMM24, Appendix B], we observe that the pentagon relation significantly simplifies the analysis (just in two lines in (5.5)).
@@ 5.3. Nonpositivity result
Consider a cluster algebra of rank 2 such that its exchange matrix $B$ has the form in (5.4). As stated in Example 5.8, it was conjectured in [LLRZ14, Conj. 14] that the positivity of the quantum greedy basis holds if $\delta_{1}|\delta_{2}$ or $\delta_{2}|\delta_{1}$. The following result is motivated by the converse of this conjecture.
@@
Let us show that under the condition the nonpositivity already emerges at degree 2 just as above Examples (a) and (d). We continue to assume that $\{e_{2},e_{1}\}=1$. By a similar consideration to the examples in Section 5.2, we have, modulo $G^{>2}$,
@@
The above result implies the nonpositivity of (quantum) theta functions.
Proposition 5.11.
Under the same condition of Theorem 5.10, there is a (quantum) theta function such that at least one of its coefficients is negative in $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$.
Proof.
Consider the theta function $\theta_{Q,\tilde{m}_{0}}$, where $Q$ is in the first quadrant and $m_{0}=(\delta_{1},-\delta_{2}-1)$. Then, there is a wall $\mathbf{w}$ in $\mathfrak{D}_{\mathfrak{s}}^{s}$ whose wall element is a negative power of $\Psi_{1/\delta((1,1)),b}[(1,1)]$. Then, the broken line $\gamma$ bending only at $\mathbf{w}$ contributing to $\theta_{Q,\tilde{m}_{0}}$ with a negative coefficient in $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$. See Figure 3. Moreover, the contribution of $\gamma$ is not canceled by any other broken lines. Indeed, such broken lines should bend exactly once both at $\mathbf{e}_{1}^{\perp}$ and $\mathbf{e}_{2}^{\perp}$, but this is impossible as seen in the figure. ∎
$Q$$\gamma$
Figure 3. Broken line for $\theta_{Q,\tilde{m}_{0}}$ in Proposition 5.11 for $(\delta_{1},\delta_{2})=(3,4)$. The dashed broken line fails to reach $Q$.
We have the following corollary of Theorem 5.10.
@@
Let $\mathfrak{D}^{q}_{\mathfrak{s}}$ be a QCSD of any rank, and let $B$ be the initial exchange matrix in (2.1). Then, $\mathfrak{D}^{q}_{\mathfrak{s}}$ is nonpositive if there is a pair $i\neq j$ with $b_{ij}\neq 0$ such that $\delta_{i}\nmid\delta_{j}$ and $\delta_{j}\nmid\delta_{i}$.
@@
Suppose that $\delta_{i}\nmid\delta_{j}$ holds for a pair $i\neq j$ with $b_{ij}\neq 0$. We consider a QCSD whose wall elements are in the form (4.21). Then, the proof of Theorem 5.10 tells that there is some wall whose wall element is a negative power of $\Psi_{1/\delta(e_{i}+e_{j}),b}[e_{i}+e_{j}]$. Moreover, due to the construction of a CSD in [GHKK18, Appendix C.1], other wall elements never cancel the above wall element. ∎
@@ 5.4. Positivity up to degree 4
We are left with the problem to determine whether the converses of Theorem 5.10 and Corollary 5.12 hold or not. Here, we concentrate on the rank 2 case and examine the positivity up to degree 4.
@@
$\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{1,0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}2\\ 3\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\\ &\qquad\times\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},-\frac{6}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},-\frac{4}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},\frac{4}{5}}\begin{bmatrix}2\\ 4\end{bmatrix}_{\frac{2}{5},\frac{6}{5}}\\ &\qquad\times\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{5},-\frac{2}{5}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{5},\frac{2}{5}}\begin{bmatrix}1\\ 4\end{bmatrix}_{\frac{1}{5},0}\begin{bmatrix}1\\ 5\end{bmatrix}_{1,0}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{5},0}.\end{split}$
@@
$\begin{split}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}&\equiv\begin{bmatrix}1\\ 0\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}2\\ 1\end{bmatrix}_{\frac{1}{4},0}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},-\frac{1}{2}}\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{4},\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{3}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-1}\\ &\qquad\times\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},1}\begin{bmatrix}2\\ 2\end{bmatrix}_{\frac{1}{2},\frac{3}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-1}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},-\frac{1}{2}}\\ &\qquad\times\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},0}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},\frac{1}{2}}\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{2},1}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},-\frac{1}{2}}\begin{bmatrix}1\\ 3\end{bmatrix}_{\frac{1}{4},\frac{1}{2}}\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{4},0}.\end{split}$
@@
(c). Degree 4. The results are similar to the degree 3 case. However, the formulas become lengthy, so we put them in Appendix B. They also demonstrate the complexity of proving the positivity in higher degrees by direct computation.
@@ Appendix B Positivity at degree 4
This is a continuation of Example 5.14 for degree 4. The purpose of writing down the explicit expressions is to demonstrate how the problem becomes complicated even for such a small degree. We have